OpenStudy (madmerc):

Simplify (x^2/3)^4/5

OpenStudy (alexandervonhumboldt2):

use the property: (a^b)^c=a^(bc). (x^2/3)^4/5=x^(2/3 * 4/5)=simplify

OpenStudy (anonymous):

When an exponent is raised to another power, multiply the exponents. (x^2/3)^4/5 x^((2/3)(4/5)) x^(8/15)

OpenStudy (madmerc):

thank you!

OpenStudy (anonymous):

Am I right @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yeah